OpenStudy (anonymous):

Solve: `3/4y - 2 = 6 - 1/4y` A. x = 4 B. x = 16 C. x = 2 D. x = 8

OpenStudy (shamil98):

there are two variables.. and the question has y but the choices have x?

OpenStudy (shamil98):

I assume you're asking to solve for x... 3/4 x - 2 = 6 - 1/4 x x = 8